Calculate the magnitude of the faintest object that a 100m optical telescope can detect?
1
Expert's answer
2020-02-27T10:06:14-0500

The penetrating power of the telescope is characterized by the limiting magnitude m of the faintest star that can be seen in this instrument.

If the diameter of the telescope

"D=100m=100000mm"

then

"m=2.1+5\cdot \lg{(D)}=2.1+5\cdot \lg{(100000)}=2.1+5 \cdot 5=27.1^m"